A company uses 2 machines to produce rods. The length of the rods from both machines follow exponential distributions. Machine A produces rods with a mean length of 6 inches, while machine b produces rod with mean length of 10 inches. In the factory 1/3 of rods are produce by machine A and 2/3 by machine B. 2 rods are selected from a defective batch, both have a length greater than 12. What is the probability the defective rods were made by machine B?
Answer #1

P(both are greater than 12) =P(from machine A and both are greater than 12)+P(from machine B and both are greater than 12)

=(1/3)*(e-12/6)2+(2/3)*(e-12/10)2=0.066584

therefore P(from machine B given  both have a length greater than 12. )

=(2/3)*(e-12/10)2/0.066584 =0.908308